A particle moves in the xy-plane with the position given by the parametric equations:


x = t^3-t^2-9t +9

y = 2-t


Find the location of the particle when = 1.


(_,_)


1
Expert's answer
2021-03-29T07:25:58-0400

A particle moves in the xy-plane with the position given by the parametric equations:

x(t) = t3 - t2 - 9t + 9

y(t) = 2 - t

we have to find the location of the particle when = 1.

"\implies" x(1) = 13 - 12 - 9"\cdot"1 + 9

x(1) = 1 - 1 - 9 + 9

x(1) = 0

"\implies" y(1) = 2 - 1

y(1) = 1

hence,

Location of the particle at t = 1 is (0,1)